OpenStudy (anonymous):

Can someone tell me how you get the 78.4 in this problem. C"(x)=(1.6(x^2+49)^-1/2-1) =78.4/(x^2+49)^3/2

OpenStudy (mertsj):

\[C"(x)=\frac{1.6}{\sqrt{x^2+49}}-1\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

I just don't understand how you get the 78.4. The problem says Find the sign of C"(5.6). C"(x)=(1.6(x^2+49)^-1/2-1)' =78.4/(x^2+49)^3/2 C"(5.6)=78.4/((5.6)^2+49)^3/2

OpenStudy (mertsj):

And I cannot tell where the 78.4 came from unless I know what the original problem is. So I guess we're at a Mexican stand off as the expression goes. I can't answer your question until you answer mine and you don't want to answer mine. Sorry.

OpenStudy (anonymous):

I don't know why but 1.6x49=78.4